with a 10% increase in the price of a dairy products , the number of units supply rises from 20 to 25 . Determine the price elasticity of supply . state whether the supply for the product is elastic or inelastic .


1
Expert's answer
2021-04-30T07:37:02-0400

Percentage increase in the price of goods = 10%

Percentage change in quantity supplied is given as:

"\frac{{P_1-P_0}}{P_0}×100"

Where "P_0" is the old price and "P_1" is the new price

Therefore percentage change in quantity supplied

= "\frac{25-20}{20}×100"

=25%

E= (%change in quantity supplied)/(% change in price)

"E=\frac{0.1}{0.25}"

E= 0.4


The supply is inelastic